OpenStudy (anonymous):

solve for x and y 3x + y = 6 4x + 4y = 16

sam (.sam.):

3x + y = 6...........x4 4x + 4y = 16 ------------------------------------------------ 12x + 4y = 24 4x + 4y = 16 ============ 8x=8 x=1 ----------------------------------------------- 4x + 4y = 16 4(1) + 4y = 16 y=3